# AutoGraphAD: A novel approach using Variational Graph Autoencoders for anomalous network flow detection

### v1

@@ Abstract
Network Intrusion Detection Systems (NIDS) are essential tools for detecting network attacks and intrusions. While extensive research has explored the use of supervised Machine Learning for attack detection and characterisation, these methods require accurately labelled datasets, which are very costly to obtain. Moreover, existing public datasets have limited and/or outdated attacks, and many of them suffer from mislabelled data. To reduce the reliance on labelled data, we propose AutoGraphAD, a novel unsupervised anomaly detection approach based on a Heterogeneous Variational Graph Autoencoder. AutoGraphAD operates on heterogeneous graphs, made from connection and IP nodes that capture network activity within a time window. The model is trained using unsupervised and contrastive learning, without relying on any labelled data. The reconstruction, structural loss, and KL divergence are then weighted and combined in an anomaly score that is then used for anomaly detection. Overall, AutoGraphAD yields the same, and in some cases better, results than previous unsupervised approaches, such as Anomal-E, but without requiring costly downstream anomaly detectors. As a result, AutoGraphAD achieves around 1.18 orders of magnitude faster training and 1.03 orders of magnitude faster inference, which represents a significant advantage for operational deployment.
1 Introduction
@@
A substantial body of research has explored the use of Machine Learning (ML) to detect and characterise attacks [8171733, LightweightIDS, 10127959, 10.1007/978-981-13-1921-1_35]. However, most of these approaches are susceptible to adversarial attacks, and therefore do not suit real-life deployment [7467366]. Graph Neural Networks are known to be more robust against adversarial attacks, since they do not rely solely on network data, but also learn the structures and relationships among network traffic flows [pujolperich2021unveilingpotentialgraphneural, 10123384, 10.1145/3664476.3664515].
In the NIDS domain, the number of datasets that can be used to create and train ML models remains limited [goldschmidt2025networkintrusiondatasetssurvey]. Most of these datasets use outdated attacks and are highly synthetic, creating a distorted image of how attacks are structured and work [PINTO2025111177].
@@
Anomaly detection is a good approach to mitigate reliance on labelled datasets, as we only need normal data to establish a baseline, and anything that deviates from it is considered anomalous. In the domain of graph-based anomaly detection for NIDS, Anomal-E is arguably the most representative and influential approach [Caville_2022]. Anomal-E constructs a homogeneous graph, where network flows are represented as edge features, which are then transformed to edge embeddings for anomaly detection. However, Anomal-E requires the generation of large embeddings, introducing significant overhead and slowing down the detection process [doi:10.1126/science.153.3731.34]. Furthermore, in the case of concept drift [Webb2016-vo], the entire pipeline has to be re-trained, making it computationally expensive to deploy in operational environments.
@@
In the case of Anomal-E, and many other approaches [arganids, zoubir2024integratinggraphneuralnetworks, 10.1007/978-3-030-50423-6_12], they rely on traditional downstream anomaly detection algorithms, which are often not suitable for a streaming environment [app13106353], which makes them very difficult to deploy in real-world networks.
In this paper, we propose AutogGraphAD, a novel approach for unsupervised anomaly detection based on Variational Graph Autoencoders (VGAEs). In our approach, we represent network traffic as a heterogeneous graph with two distinct types of nodes, IP and Connection nodes. We then performed unsupervised contrastive training for the VGAE, without requiring any labelled data. We leverage the generated embeddings to calculate anomaly scores that are then used for anomaly detection. We have tested multiple different architectural variations to determine which implementation yields the best results at different levels of contamination (i.e., presence of anomalous flows) of the training dataset. AutoGraphAD yields the same, and in some cases better, results than Anomal-E while not relying on the use of costly downstream estimators. Thus, faster training by 1.18 orders of magnitude and faster inference by 1.03 orders of magnitude is achieved, showcasing a significant advantage for operational deployment.
2 Problem Description
As mentioned in Section 1, the main problem that this paper aims to solve is the detection of network attacks without the need for expensive and limited labelled data. This is a very important problem that needs to be solved if we want to achieve more efficient results and a better real-life deployment of Machine Learning based NIDS.
The main issue with current solutions such as Anomal-E [Caville_2022] and ARGANIDS [arganids] is the reliance on downstream detectors such as HBOS [goldstein2012hbos] to classify the embeddings generated for classification purposes. This adds an additional layer of computation that needs to be retuned often to account for concept drift [Webb2016-vo] and the network’s dynamic nature [app13106353].
@@ 2 Problem Description
Moreover, certain one class classification algorithms tend to become prohibitively expensive when depending on the size of the embedding used. This can be seen in the case of OneClassSVM [NIPS1999_8725fb77] and its linear variants, where the dimensionality of the features can negatively impact training and inference time, as well as memory complexity and overall performance [scikit-learn]. It should be noted that, in the case of Anomal-E [Caville_2022], the reported results use 256 vector embeddings. These embeddings can be considered large and affect anomaly detection algorithms as they suffer from the curse of dimensionality [doi:10.1126/science.153.3731.34].
@@
An interesting approach to avoid reliance on a downstream estimator is the use of dynamic Graph Autoencoders (VGAEs) to detect anomalies. This approach has been used successfully in network monitoring as seen in GAT-AD [LATIFMARTINEZ2025110830]. In the case of GAT-AD, the predicted and actual values are used to calculate an anomaly score, and then, based on those node scores, we can detect which nodes are anomalous and which are not.
In the case of anomaly detection in a network setting, the majority of state-of-the-art approaches rely on downstream estimators to conduct the anomaly detection.
In this paper, our aim is to achieve network detection without relying on a downstream detector. We rely solely on reconstruction errors and metrics and not on any other downstream detectors. We are aiming to leverage VGAEs and their reconstruction errors as a way to detect anomalous nodes. AutoGraphAD is substantially faster computationally than Anomal-E’s and ARGANID’s approaches, while much simpler and much more easily tunable.
3 Background
In this section, we describe the main pre-requisites of the technologies that we will be using in this paper as part of developing our solution.
3.1 Variational Graph Autoencoders
The underlying technology that we will use in this paper is VGAEs. VGAEs were originally proposed by Kipf in 2016 [kipf2016variationalgraphautoencoders].
They are a Graph Neural Network extension of traditional Variational Autoencoders [kingma2022autoencodingvariationalbayes]. Their initial design relied on the reconstruction of the graph structure combined with the loss of Kullback-Leibler regularisation [kipf2016variationalgraphautoencoders].
The variational autoencoder works by the encoder component creating uncertainty-aware node embeddings that are then parametrised into node embeddings and coupled with a Dot Product to recreate the Adjacency matrix of the graph [kipf2016variationalgraphautoencoders]. Then both adjacency matrices are used to compute the reconstruction error. The equations used to calculate the errors can be seen in the following: Equation 1 is the Dot Product, Equation 2 is the Binary Cross Entropy Loss used to compute the structural reconstruction error [kipf2016variationalgraphautoencoders], of Equation 3 is the KL regularisation error, and Equation 4 is the total loss used.
$A^{\prime}=Z\cdot Z^{T}$ (1)
$L_{Recon}=BCE(A^{\prime},A)$ (2)
$KL_{Reg}=-\frac{1}{2}*\sum(1+\log(\sigma^{2})-\mu^{2}-\sigma^{2})$ (3)
$L_{Total}=L_{Rrecon}+KL_{Reg}$ (4)
3.2 GraphSAGE
GraphSAGE is a Graph Neural Network architecture that can learn to generate embeddings from sampling and aggregating the node’s neighbours [NIPS2017_5dd9db5e]. After the aggregation, the generated embedding goes through a non-linear update.
The main advantage of GraphSAGE compared to other algorithms is that it can easily work with unseen nodes [NIPS2017_5dd9db5e], making it ideal for real-time deployment where the nodes and graph structure differ from training data. In addition, GraphSAGE has been proven to scale very well in larger graphs [NIPS2017_5dd9db5e], making it an ideal choice for a large network deployment.
3.3 Adversarially Regularised VGAEs
Adversarially Regularised VGAEs [pan2019adversariallyregularizedgraphautoencoder] is another approach to train the autoencoders to generate higher quality embeddings. This approach uses the exact same approach as the original VGAE paper [kipf2016variationalgraphautoencoders] but adds the use of a regulariser model to improve the performance. The regulariser works by taking samples from the generated embeddings and from the a Gaussian distribution and learns to separate which one is which. This is then fed into the total loss that is used for back-propagation.
3.4 Multiple Sampling Variational Autoencoders
Another upgrade that can be done to enhance the VGAEs is through the use of multiple latent embeddings per node. The idea behind using multiple embeddings that are then used to reconstruct multiple reconstruction losses. This allows for the generation of an unbiased estimate of the same objective with a much smaller variance [pmlr-v48-mnihb16]. This has been proven to improve the performance of the model and yield better results [pmlr-v80-buchholz18a].
3.5 KL Annealing
KL Annealing can be used to improve the performance of VAEs [bowman2016generating]. KL Annealing works by slowly increasing the KL regularisation loss through a set number of epochs. By slowly increasing the effect of the KL loss, the autoencoder is initially forced to learn meaningful embeddings, and then transitions on enforcing prior matching. The goal of this approach is to mitigate the posterior collapse, where KL transitions to 0, which can negatively affect the performance of the VAE.
3.6 Graph Masked Autoencoders
Masking was proposed by Hou et al. [GraphMAE] in the Graph Masked Autoencoder (GraphMAE) and used in the Heterogenous Graph Masked Autoencoder (HGMAE) [tian2023heterogeneousgraphmaskedautoencoders] to enhance the performance of their model. Their GAE architecture is enhanced by masking the input node features as well as the latent embeddings. Additionally, as part of model training, the graph edges are dropped. For masking, they use a learnable masking token similar to the one used in BERT [devlin2019bertpretrainingdeepbidirectional]. Masking and edge dropping make it harder for the autoencoder to reconstruct the information, and thus the autoencoder does not learn trivial solutions during training but rather focusses on learning the node relations.
4 Methodology
@@
In this section, we focus on describing our methodology and how we have created, trained, and evaluated AutoGraphAD. Additionally, we describe our pre-processing, graph-building, and dataset creation pipelines.
4.1 Dataset Used
@@
For this work, we have chosen to use the UNSW-NB15 dataset [7348942]. We chose UNSW-NB15 for the AutoGraphAD evaluation, as it is a widely used dataset with clearly labelled modern attacks. Thus, our evaluation is better suited for the current environment. The data set has been generated from the raw PCAP files using NFStream [AOUINI2022108719]. For the generation of the flows, we used the same settings as were used in the original dataset generation. NFStream can output flows with multiple features. We chose the features that overlap with the features present in NetFlow V9 [claise2004cisco].
For UNSW-NB15, we used the generated labels from the original paper [7348942] and matched them with the flows generated by NFStream. The features used can be found in Table V.
4.2 Dataset Creation
@@
Having described the data pre-processing, it is important to describe how the datasets are created. The goal of this paper is to explore and showcase our model’s ability to detect anomalies even when the training data are contaminated. To achieve this, we have created different training parts with different levels of contamination similar to Anomal-E [Caville_2022]. The main difference between Anomal-E’s approach and ours is that Anomal-E uses 10% of the data set to create a single large graph for the training, and the same is repeated for testing. Our goal is to create 3 levels of contamination: No contamination, 3.5% Contamination, and 5.7% Contamination.
In our case, we use time windows to detect attacks and simulate a real-life deployment. For the training split of the dataset, we follow the following procedure; once we create all the sliding windows for the entire dataset, we calculate if the windows are anomalous or not, meaning if they contain attacks or not, and we do a stratified splitting to all datasets. For a training dataset that does not have any attacks, we removed all the anomalous graphs. For 3.5% contamination, we keep the stratified split, and for 5.7% we downsample the benign graphs for UNSW-NB15 as the normal contamination of the level is around 3.5%.
4.3 Feature Pre-processing and Graph Creation
For feature pre-processing, we use a similar pre-processing approach to the one proposed in Anomal-E [Caville_2022]. The dataset is initially checked for flows with missing or infinite values, which are then removed. Following this step, we begin by building the time windows that will be used for training, testing, and evaluation. We then split the randomly generated graphs into parts for training, evaluation, and testing. We allocate 70% of the data for training, 20% for evaluation, and 10% for testing. For the time windows, we chose to create non-overlapping graphs that capture 180 seconds of traffic.
The second step of the pre-processing is feature normalisation. For categorical features, which in our case is the IP version and protocols, we employ One-Hot encoding [Hancock2020]. In the case of the protocols, we use Rare Labelling from the library Feature Engine [Galli2021] due to the incredibly large number of protocols and the fact that the majority of these are very sparse in the data. Including these features in One-Hot encoding would increase the dimensionality of the data without yielding any performance gain; as such, the rarer protocols are bundled in a third category labelled Rare, resulting in three One-Hot categories, two for the two most common features, and one for all the rare features.
For the continuous features, we are using L2 row wise normalisation. This choice was made because it does not require any significant pre-processing and can be done on the fly in a real-life environment, allowing for an easier real-life deployment. The entire process from dataset creation to graph creation can be seen in Figure 1.
Figure 1: Dataset Pre-processing pipeline for graph generation
4.4 Network Architecture and Variants
@@
Figure 2: The proposed Architecture of AutoGraphAD in a training setting. AutoGraphAD mainly focuses on the reconstruction and use of the connection nodes as the IP nodes have placeholder values. The pipeline starts with the encoder that generates the latent space embeddings that are then reparametised to create the embeddings that are then used for structure and feature reconstruction. Using the reconstructed values, we calculate the reconstruction losses used for back-propagation. The GNN encoder and decoder can be switched out with every other GNN algorithm. The same can be done regarding the losses that are used in Back-propagation.
@@
Our proposed architecture builds upon the original VGAE [kipf2016variationalgraphautoencoders] and extends it by incorporating elements from GraphMAE [GraphMAE] and HGMAE [tian2023heterogeneousgraphmaskedautoencoders]. Our architecture is designed to reconstruct both node features and graph structure during training. The main upgrades introduced in our design are as follows.
@@
We are using the same masking component as used in GraphMAE [GraphMAE] and HGMAE [tian2023heterogeneousgraphmaskedautoencoders], but instead of masking both latent embeddings and original graph embeddings, we only mask the original graph nodes. We made this decision because during the evaluation of AutoGraphAD there was a negligible improvement when using double masking.
@@
We incorporated random edge drop during training to force the model to learn non-trivial relationships between nodes [tian2023heterogeneousgraphmaskedautoencoders]. We have also incorporated negative edge sampling as a contrastive learning objective to help the autoencoder to better distinguish between real and fake edges. This helps our autoencoder to learn better the structure of the graphs.
@@
Regarding feature reconstruction, a GNN has been employed to decode the node features. A GNN was chosen as it is the most capable of decoding the embeddings correctly. GNNs as a decoder are used in both HGMAE and GraphMAE [tian2023heterogeneousgraphmaskedautoencoders, GraphMAE], but in these approaches they have focused only on decoding masked embeddings. In this architecture, all nodes are decoded. In our approach, we only decode the connection nodes as the IP nodes are placeholder values that do not contain any meaningful information.
Regarding the feature reconstruction errors, we have created two main variants, one using the Cosine Embedding Error and one using the Mean Squared Error (MSE). The Cosine Embedding Error approach has been inspired by HGMAE [tian2023heterogeneousgraphmaskedautoencoders] and GraphMAE [GraphMAE], while the MSE approach has been chosen as the goal here of the autoencoder is to be able to generate meaningful embeddings in which the decoded values must be closest to the actual values. The KL divergence loss used in our model is the same as in the original VGAE paper [kipf2016variationalgraphautoencoders].
For the total reconstruction loss, we combine the feature reconstruction loss with the structural reconstruction loss and KL divergence. The total loss is then back-propagated through the model to update the model. Each component of the total reconstruction loss, with the exception of the KL divergence, is weighted by a user defined value. The model architecture can be found in Figure 2 and the total loss is depicted in equation 6.
$L_{Total}=\alpha*L_{Structure}+\beta*L_{Features}+KL$ (6)
4.5 Model Training
The model was trained with 100 maximum allowable epochs. We used an early stopping mechanism during the training to prevent overfitting and expenditure of the computational resources. Furthermore, we saved the model’s weights at the point that the model achieved its best performance. We ensured that we maintain the optimal version of the model.
4.6 Anomaly Detection
For anomaly detection, we are calculating the anomaly scores of the connection nodes. The anomaly scores are based on the calculation of the reconstruction errors with the KL term. The idea is that the anomaly score will be lower for normal nodes and higher for anomalous nodes. To determine the threshold of what an anomaly is and what it is not, we use percentiles to determine a user defined percentile, and then everything above it is determined as anomalous scores. For our score, we use the following equation.

### v2

# AutoGraphAD: Unsupervised network anomaly detection using Variational Graph Autoencoders

@@ Abstract
Network Intrusion Detection Systems (NIDS) are essential tools for detecting network attacks and intrusions. While extensive research has explored the use of supervised Machine Learning for attack detection and characterisation, these methods require accurately labelled datasets, which are very costly to obtain. Moreover, existing public datasets have limited and/or outdated attacks, and many of them suffer from mislabelled data. To reduce the reliance on labelled data, we propose AutoGraphAD, a novel unsupervised anomaly detection based on a Heterogeneous Variational Graph Autoencoder. AutoGraphAD operates on heterogeneous graphs, made from connection and IP nodes that represent network activity. The model is trained using unsupervised and contrastive learning, without relying on any labelled data. The model’s losses are then weighted and combined in an anomaly score used for anomaly detection. Overall, AutoGraphAD yields the same, and in some cases better, results than Anomal-E, but without requiring costly downstream anomaly detectors. As a result, AutoGraphAD achieves around 1.18 orders of magnitude faster training and 1.03 orders of magnitude faster inference, which represents a significant advantage for operational deployment.
I Introduction
@@
A substantial body of research has explored the use of Machine Learning (ML) to detect and characterise attacks [8171733, LightweightIDS, 10127959, 10.1007/978-981-13-1921-1_35]. However, most of these approaches are susceptible to adversarial attacks and therefore do not suit real-life deployment [7467366]. Graph Neural Networks (GNNs) are known to be more robust against adversarial attacks, as they use both network data and learn the structures and relationships among network traffic flows [pujolperich2021unveilingpotentialgraphneural, 10123384, 10.1145/3664476.3664515].
In the NIDS domain, the number of datasets that can be used to create and train ML models remains limited [goldschmidt2025networkintrusiondatasetssurvey]. Many public NIDS datasets contain synthetic elements, outdated attacks, or labelling limitations, creating a distorted image of how attacks are structured and work [PINTO2025111177].
@@
Anomaly detection is a good approach to mitigate reliance on labelled datasets, as we only need normal data to establish a baseline, and anything that deviates from it is considered anomalous. Anomal-E is arguably the most representative and influential approach in graph-based network anomaly detection [Caville_2022].
@@
In the case of Anomal-E and other approaches [zoubir2024integratinggraphneuralnetworks, 10.1007/978-3-030-50423-6_12], they rely on traditional downstream anomaly detection algorithms, which are often not suitable for a streaming environment [app13106353], making them very difficult to implement in real-world networks. This is an additional layer of computation that needs to be retuned often to account for concept drift [Webb2016-vo] and the network’s dynamic nature [app13106353]. As both the model and the estimator would need to be taken offline to retrain and then redeploy.
@@
Moreover, certain one class classification algorithms tend to become prohibitively expensive when depending on the size of the embedding used. This can be seen in the case of OneClassSVM [NIPS1999_8725fb77] and its linear variants, where the dimensionality of the features negatively affects training and inference time, memory complexity, and overall performance [scikit-learn]. Anomal-E [Caville_2022] uses 256 vector embeddings, which are large and can negatively affect detection algorithms.
@@
An interesting approach to avoid reliance on a downstream estimator is the use of dynamic Graph Autoencoders (GAEs) to detect anomalies. This approach has been used successfully in network monitoring as seen in GAT-AD [LATIFMARTINEZ2025110830], where the deviation of the predicted values from the actual values is used to detect anomalies.
In this paper, we propose AutoGraphAD, a novel approach to unsupervised anomaly detection based on Variational Graph Autoencoders (VGAEs). In our approach, we represent network traffic as a heterogeneous graph with two distinct types of nodes, IP and Connection nodes. We then performed unsupervised contrastive training for the VGAE, without requiring any labelled data. We rely solely on reconstruction errors and metrics to calculate the anomaly score, and we do not rely on any other downstream detectors. We have tested multiple different architectural variations to determine which implementation yields the best results at different levels of contamination (i.e., presence of anomalous flows) of the training dataset. The choice of testing at different contamination levels arises as we want to evaluate how robust the proposed architecture is. AutoGraphAD yields the same, and in some cases better, results than Anomal-E while not relying on the use of costly downstream estimators. Thus, faster training by 1.18 orders of magnitude and faster inference by 1.03 orders of magnitude is achieved, showcasing a significant advantage for operational deployment. We evaluate the method under multiple contamination levels and show comparable detection quality to Anomal-E while substantially reducing training and inference time.
II Related Work
NIDS are a broad area with thousands of proposals; a review of all proposals is outside the scope of this paper. Thus, we will focus on related work in terms of GNN-based NIDS. Surveys on NIDS and GNN-based NIDS can be found here [goldschmidt2025networkintrusiondatasetssurvey, ZHONG2024103821].
GNN-based NIDS are mostly supervised. E-GraphSAGE [Lo_2022] is a widely used representative baseline as it is a widely evaluated flow-graph GNN [Venturi2023]. The model works by implementing E-GraphSAGE, a variation of GraphSAGE that works by processing only the Edge features rather than the node features that traditional GraphSAGE captures [NIPS2017_5dd9db5e]. It is trained to perform edge (flow) classification. GraphIDS is a recently published model based on E-GraphSAGE [guerra2025selfsupervisedlearninggraphrepresentations] by augmenting flow representation learning with a masked encoder-decoder Transformer. ARGANIDS [arganids] uses Adversarially Regularised Graph Autoencoder (ARGA) [pan2019adversariallyregularizedgraphautoencoder] GAE to learn node embeddings and then uses a downstream supervised Machine Learning classifier.
As mentioned in the Introduction, the main limitation of using supervised learning is the reliance on limited and possibly outdated datasets [PINTO2025111177] due to the increased cost of labelling network data [Braun_2024]. This cannot easily scale for a real-life implementation, leading to the need for unsupervised approaches.
When it comes to unsupervised approaches, there is much less related work. Anomal-E is a state of the art network anomaly detection algorithm that implements a GAE to learn meaningful embeddings of a graph [Caville_2022]. Anomal-E implements E-GraphSAGE. A modified Graph Infomax version is then used to train the model encoder. The resulting embeddings are then fed to a downstream estimator used for anomaly detection. Although this approach is effective, the pipeline adds more complexity through additional components and tuning options. As such, we compare primarily unsupervised approaches, as we assume that there are no labels available during training.
III Methodology
@@
In this section, we first describe the dataset construction and graph representation. We then focus on describing our methodology and how we have created, trained, and evaluated AutoGraphAD. Finally, we explain the anomaly scoring and evaluation.
III-A Dataset
@@
For this work, we used the UNSW-NB15 dataset [7348942]. We chose UNSW-NB15 for the AutoGraphAD evaluation, as it is a widely used dataset with clearly labelled modern attacks. Thus, our evaluation is better suited for the current environment. The data set has been generated from the raw PCAP files using NFStream [AOUINI2022108719], which outputs flows with NetFlow V9 compatible features [claise2004cisco]. To create the labels, we used the UNSW-NB15 generated labels [7348942] and matched them to the generated flows.
@@
The goal of this paper is to explore and showcase our model’s ability to detect anomalies even when the training data are contaminated. To achieve this, we have created different training parts with different levels of contamination similar to Anomal-E [Caville_2022]. The main difference between Anomal-E’s approach and ours is that Anomal-E uses 10% of the data set to create a single large graph for training, and the same is repeated for testing.
We use time windows to detect attacks and simulate real-life deployments. To achieve a fair distribution of attacks, we use stratified data split to see whether or not contaminated attacks are equally distributed. For 0% contamination, we remove all the graphs with attacks, for contamination, we use a stratified split of the dataset, and for 5.7% we downsample the benign graphs to of their original number.
We allocate 70% of the data for training, 20% for evaluation, and 10% for testing. For the time windows, we chose to create non-overlapping graphs that capture 180 seconds of traffic. For categorical features, which in our case is the IP version and protocols, we employ One-Hot encoding. In the case of the protocols, we use Rare Labelling from the library Feature Engine [Galli2021] to reduce the dimensionality of the data. For continuous features, we use L2 row-wise normalisation.
The graph is made up of IP nodes that contain placeholder values used for message passing and connection nodes that contain flow features. The graph is undirected.
@@
Figure 1: The proposed Architecture of AutoGraphAD in a training setting. AutoGraphAD mainly focuses on the reconstruction and use of the connection nodes as the IP nodes have placeholder values. The pipeline starts with the encoder that generates the latent space embeddings that are then reparametised to create the embeddings that are then used for structure and feature reconstruction. The GNN encoder and decoder as well as the losses can be switched out with every other GNN algorithm. In this paper, we are employing GraphSAGE as the encoder and decoder algorithm.
III-B Model Architecture
@@
Our proposed architecture builds on the original VGAE [kipf2016variationalgraphautoencoders] and extends it by incorporating elements from GraphMAE [GraphMAE] and HGMAE [tian2023heterogeneousgraphmaskedautoencoders]. Our architecture is designed to reconstruct both the node features and the graph structure during training. The main upgrades introduced in our design are as follows.
@@
We are using the same masking component as used in GraphMAE [GraphMAE] and HGMAE [tian2023heterogeneousgraphmaskedautoencoders], but we only mask the original graph nodes. We made this decision because during the evaluation of AutoGraphAD there was a negligible improvement when using double masking.
@@
We incorporated random edge drop during training to force the model to learn non-trivial relationships between nodes [tian2023heterogeneousgraphmaskedautoencoders]. We have also incorporated negative edge sampling as a contrastive learning objective to help the autoencoder to better distinguish between real and fake edges.
@@
Regarding feature reconstruction, a GNN model has been employed to decode only the connection node features. A GNN was chosen as it is the most capable of decoding the embeddings correctly. GNNs as a decoder are used in both HGMAE and GraphMAE [tian2023heterogeneousgraphmaskedautoencoders, GraphMAE], but in these approaches they have focused only on decoding masked embeddings. In this architecture, all of the connection nodes are decoded.
Regarding the feature reconstruction errors, we have created two main variants, one using the Cosine Embedding Error and one using the Mean Squared Error (MSE). The Cosine Embedding Error approach has been inspired by HGMAE [tian2023heterogeneousgraphmaskedautoencoders] and GraphMAE [GraphMAE], while the MSE approach was chosen to help generate meaningful embeddings that are closest to the actual values. The KL divergence loss used in our model is the same as in the original VGAE paper [kipf2016variationalgraphautoencoders].
For the total reconstruction loss, we combine the feature reconstruction loss with the structural reconstruction loss and KL divergence. Each component of the total reconstruction loss, with the exception of the KL divergence, is weighted by a user-defined value. The model architecture can be found in Figure 1 and the total loss is depicted in Equation 2.
$L_{Total}=\alpha*L_{Struct}+\beta*L_{Feat}+KL$ (2)
The model was trained with 100 maximum allowable epochs. We used an early stopping mechanism during the training to prevent overfitting and expenditure of the computational resources. Furthermore, we saved the model’s weights at the point that the model achieved its best performance.
III-C Anomaly Detection
For anomaly detection, we calculate the anomaly scores of the connection nodes. The anomaly scores are based on the calculation of the reconstruction errors with the KL term. The idea is that the anomaly score will be lower for normal nodes and higher for anomalous nodes. We use a percentile-based threshold, where a node is classified as anomalous if its score exceeds the p-th percentile; $p$ is selected offline as a threshold hyperparameter. For our score, we use the following equation.